# WakenLLM: Evaluating Reasoning Potential and Stability in LLMs via Fine-Grained Benchmarking

### v4

@@ 3 Experiments
3.1 Benchmarks
To explore the effects of the Vague Perception phenomenon on different types of logical datasets, we classify them based on their labels: (1) Verifiable Data, where the samples have a clear True/False/Unknown label indicating the validity of the reasoning; and (2) Unverifiable Data, where the label is Unknown and samples are unsolvable. Furthermore, as shown in Figure 3, to better understand how LLMs reason across different logical forms, we categorize benchmarks as follows:
Fact-based Benchmarks.
The information is given in atomic relationships, split by unrelated sentences.: FLD (Morishita et al., 2024) supplies three kinds of sample labels annotated as True, False, or Unknown. We randomly sample 600 items: 300 verifiable (True/False) and 300 marked Unknown due to insufficient evidence. FOLIO (Han et al., 2022) is a factual benchmark with the same 3-way labeling. Similarly, we draw 640 samples—320 verifiable samples with T/F labels and 320 Unknown.
Story-based Benchmarks.
Our experimental design considers the differing characteristics of the two forms above. For Fact-based ones, which models tend to process the entire context holistically, so providing additional stimulation can enhance performance. In contrast, we extract samples from ScienceQA (Lu et al., 2022) for story-based tasks—often rooted in professional knowledge domains and relying on models’ inherent knowledge—we minimize guidance to avoid interfering with the model’s reasoning process.
@@ 3.2 Experimental Setup
Our evaluation covers both closed-source models: ChatGPT-3.5-turbo and GPT-4o (OpenAI, 2023), and open-source models: LLaMA-3.1 series (Touvron et al., 2023), and Qwen-2.5 series (Bai et al., 2023). Two stimulating strategies are implemented:
@@
$\triangleright$ Concise Stimulation: Basic error notification without correction guidance.
$\triangleright$ Detailed Stimulation: Enhanced detailed approach featuring:(1) Listed Conditions Guidance.(2) Structured Reasoning Constraints. (3) Conditional Verification Steps.The experimental protocol differs between approaches. Concise Stimulation uses a similar format for two-stage error notification process, while Detailed Stimulation employs a different two-stage workflow: a standard prediction followed by an interrogative refinement. The latter stimulates analytical reasoning through questioning and step-by-step error analysis, creating a clear methodological contrast.
@@
To achieve a balance between randomness and coherence, we set all temperature to 0.5, except for 1.0 in ablation study where we want to further examine the quantitative influence between prompt designing and models setup.
In our result detection procedure, the output is first examined for specific words. When such methods are either absent or multiple candidates are detected, GPT-4o is employed to adjudicate and assign the appropriate label. With the exception of the Remind-then-Guide (RtG) setting in the reasoning process, most configurations yield a single label accompanied by few explanations, which facilitates reliable identification.
3.3 Results on Benchmark Converting Rate
@@
As shown in Table 3, we report the overall benchmark Converting Rate, which means three metrics in three systems: True Converting Rate (TCR) for both stages, and Overall Converting Rate (OCR) taking the portion of Vague Perception samples into consideration. The LLaMA and Qwen series show superior Overall Converting Rate (OCR) on small-parameter series, whereas the GPT series shows the opposite trend. We infer that, in general, higher parameter counts correlate with lower OCR. This situation may change due to the distinct yet controversial architecture of GPT-4o.
Table 4: Different Stimulating Method Influence on Converting Rate, where $\mathrm{TCR}^{1}=\mathrm{TCR}_{u}^{1}+\mathrm{TCR}_{v}^{1}$ as we split converting rate by verifiable and unverifiable samples. CS, DS refer to Concise and Detailed Stimulation. Rows shaded gray denote ones used in our framework.
Metrics(%) TCR1 $_{v}^{1} TCR$$_{u}^{1}$ Reliability
@@
3.4 Results on RtG Conformity Stimulating
As illustrated in the Figure 2, RtG module contains two parts: Label and Reasoning Process (RP) Conformity: We remind the model incorrectly predicted the sample as Unknown before; we then we apply conformity test to verify whether models can inference independently.
Label Conformity
Judging from the table, we observe that even when LLMs are reminded of their previous incorrect answers, they still demonstrate a strong tendency toward conformity. Moreover, distinct conformity patterns emerge between verifiable and unverifiable samples. One key influencing factor appears to be the model size: models with smaller parameter sizes, such as GPT-3.5 and LLaMA-3.1-8B-Instruct, are more inclined to being misguided by verifiable samples with T/F labels reasoning. In contrast, models with larger parameter sizes, such as GPT-4o and LLaMA-3.1-405B, tend to be more easily misguided regarding unverifiable samples with Unknown label. Meanwhile, Qwen series shows distinct yet related features compared with the above, we infer it is because parameters gap between Qwen 7B and 72B is much smaller than the former two series.
Reasoning Process (RP) Conformity
We input the previous reasoning process and explicitly ask the LLM to reflect on its mistakes, the resulting CGR values are mostly below zero, which indicates that even after being reminded of a past error and provided with its own flawed reasoning, the LLM still tends to repeat the same incorrect conclusion, leading to an even larger decline in accuracy.
Our results on RPC reveal that the reasoning process is often highly unstable. Changes in accuracy can largely be attributed to divergence in how the problem is interpreted. Notably, for unverifiable samples, the reasoning process tends to be unreliable — feeding it back into the model typically leads to a further drop in accuracy. This exposes a hidden issue: when LLMs automatically generate solutions for a dataset using a reasoning process, unverifiable problems are particularly susceptible to unreliability. Even for verifiable samples, the reasoning may remain unstable. As a result, the accuracy and reliability of AI-annotated problem-solving methods for datasets remain a problem.
4 Detailed Analysis
4.1 True Converting Rate (TCR) Analysis
As shown in the line chart of Figure 4, by comparing TCR1 and TCR2 across different benchmarks, we observe clear divergences: This indicates that many samples mis-predicted in Stage 1 can be self-corrected in Stage 2, where the model refines its initial output. In contrast, Story-based TCR1 shows a substantially lower TCR2. Whereas LLaMA-3.1-8B demonstrates much higher TCR1.
Thus, distinct focus strategies are recommended for multi-stage frameworks for the following LLMs: (i) For GPT-4o, when facing Fact-based benchmarks, broader exploratory freedom is acceptable in Stage 1. “making mistakes at first is acceptable", since later refinement proves beneficial; For Story-based benchmarks, the improvement of the subsequent stage is limited, so initial accuracy should be emphasized (ii) For LLaMA-3.1-8B, it is always better to “do it right the first time" .
4.2 Benchmark-wise Vague Perception
As each model differs in its reasoning ability, we first quantify the percentage of data in which models induce Vague Perception. The Bar chart in blue Fig 4 reports the percentage for GPT-4o—the most widely used LLM at present—while complete results for the remaining models in Appendix 7.
We can conclude that the Vague Perception phenomenon is prevalent across all benchmarks. For GPT-4o, FOLIO exhibits a relatively lower percentage. For FLD and Arts—which focus purely on language understanding, the model demonstrates greater uncertainty regarding explaining why certain samples are labeled as Unknown; In stark contrast, Science—which comprises Physics and Biology topics—faces significant challenges even when the samples are verifiable; For LLaMA-3.1-8B, there is a significant rise in the percentage of verifiable samples predicted as Unknown, but the ability to comprehend unverifiable samples is stronger.
[…]
4.3 Stimulating Style Comparison
[…]
Furthermore, we observe a latent trade-off: as the ability to resolve verifiable samples declines, the model becomes more confident on unverifiable ones. We attribute this to a shift in the model’s reasoning habits.
4.4 Ablation Study
From the Stage 1 output, we isolated the Unexcited samples—those that failed to even generate a complete logic chain—and fed them back through the same Stage 1 prompting protocol. Both LLaMA-3.1-8b-Instruct and GPT-4o are set with 1.0 temperature to ensure randomness.
As Figure 4 shows, this second pass produces virtually no additional conversions: more than 95% of the samples remain unchanged. This finding demonstrates that Vague Perception is largely immune to repeated, identical n-shot prompting; applying the same stimulation twice is therefore futile. Consequently, any effective correction must rely on a distinct stage 2 rather than reiterating stage 1, adjusting temperature alone without revising stimulation style can be limited.
Figure 5: Distribution of the root causes behind GPT-4o’s Vague Perception errors across all benchmarks (left). Qwen-2.5-7B and 72B model False and Unexcited Converting Rate comparison (Right). “FU”: Fact Understanding, “RG”: Reasoning Gap, and “Else” aggregates residual minor factors.
4.5 Baseline Comparison
In our work, we primarily analyze the “waken ability” with respect to Vague Perception samples by computing the proportion of truly converted samples in both stages ($TC^{1}$ and $TC^{2}$) relative to the percentage of Vague Perception samples; however, this constitutes only one component, we combine this component with the vanilla accuracy by adding them together, thus we gain an empirical bound of reasoning accuracy—namely, the maximum reasoning capability that LLMs can achieve without additional training or data-processing. The empirical accuracy of WakenLLM is defined by two parts: (1) the proportion of samples correctly reasoned by the LLM, and (2) the truly converted samples across both stages ($TC^{1}$, $TC^{2}$) that are potentially solvable. Acc(Vanilla) denotes the reasoning accuracy under direct prompt setting, without any intermediate interpretation. Here, D represents the complete dataset, whereas $D_{VP}$ only represents a shortlisted ones that induces Vague Perception.
Acc(WakenLLM) =
@@ 4.5 Baseline Comparison
$\underbrace{\text{Acc(Vanilla)}}_{\text{Explicit Ability}}\;+\;\underbrace{|TC^{1}\cup TC^{2}|/|D|}_{\text{Latent Ability}}$
As illustrated in the left section of the line chart in Figure 5, most baseline methods exhibit a substantial reasoning gap compared to our pipeline, indicating most current frameworks failed to trigger models’ full potential. Refer to the Table 6 for detailed experimental results.
@@ 4.6 Phenomenon Exploration
To probe the root causes of GPT-4o’s mistakes, we annotate each error sample with one of three labels:
$\triangleright$ Fact Understanding (FU) – The model does not correctly retrieve, align, or synthesize the relevant facts stated in the passage.
$\triangleright$ Reasoning Gap (RG) – The model retrieves the facts but fails to perform the intermediate logical operations that connect facts to the hypothesis.
$\triangleright$ Else – For the two above generally covers 80% or more scenarios. And there is a catch-all category for other less frequent factors. Notably, issues related to Excessive Caution are particularly prominent in this part.
Excessive Caution
Our experiments reveal that Excessive Caution—instances where the model responds with Unknown despite sufficient evidence—is nearly absent in Fact-based benchmarks but increases from 2.4% in Science to 5.4% in Arts. We attribute this rise to the narrative complexity of Story-based tasks: unlike the atomic and structured format of Fact-based benchmark, passages in Arts often embed literary allusions and background knowledge, heightening epistemic uncertainty.
Table 5: Performance comparison of Degradation Analysis Phenomenon between GPT‑4o and GPT‑3.5.
@@
[…]
4.7 Degradation Analysis
Within the pipeline, we don’t exclude the possibility that previous incorrectly predicted samples, FC2 from Stage 1 may be converted to Unknown, which should be denoted as UC2. As it is reasonable that at first, the model may generate an incorrect answer, but after further stimulation, due to its limited ability, it fails to arrive at the correct answer and can only output Unknown. We elaborate this as:
As shown in Table 5, only a small proportion of samples exhibit this phenomenon. However, in the case of GPT-4o on the reasoning-based Science dataset, the issue becomes more serious.
As for the Science dataset, GPT-4o achieves the highest TCR2 among all four datasets. Coupled with higher Deg value, this suggests that when confronted with problems it previously answered incorrectly—especially those beyond its capability—GPT-4o tends to respond with an Unknown label. In contrast, GPT-3.5-turbo is much more likely to provide a definitive but incorrect answer instead of admitting uncertainty.
This can be further proved by the uniform distribution of UCR and FCR in Appendix D acquired similar to formula TCR, denoting Unknown and False Converted rate in the Vague Perception input. We compare answering of Qwen-2.5 series (7B and 72B model), as shown in the radar-chart in Figure 5: (1) Models show distinctive yet uniform preferences among four different forms of dataset. For FLD and FOLIO, models are much more willing to output a false answer rather than admitting Unknown, this is even more serious for FOLIO dataset, where model likes to “show off" when facing easier ones, as FOLIO achieves the highest accuracy among above four; Story-based dataset, like Science and Arts are exactly the opposite. (2) The higher the parameters, the more likely models are to output Unknown when inputting problems beyond LLMs’ ability.

### v5

@@ 3.3 Remind-then-Guide (RtG) Setting
As indicated in Section 3.2, the subscript t represents two sample types s or u, separating conformity analysis by solvable and unsolvable samples. $TCR^{1\prime}_{t}$, $TCR^{2\prime}_{t}$ indicates stage-wise True Converting Rate given all assigned labels are wrong.
Reasoning Process (RP) Conformity
We also examine whether LLMs, after being reminded of previous mistakes, can reflect on previous reasoning processes to avoid making same mistakes. To the best of our knowledge, we are the first to study conformity on sentence-level answers.
Reasoning Process Coherence (RPC)
Including LLMs’ reasoning processes in prompts, and we rerun stage 1 to assess its stability. Similarly, RPC is calculated by changes of converting rate with and without reasoning processes. Values below zero indicates that the previous reasoning process is misguiding and thus results in declined accuracy, while above zero suggests that it is helpful and can guide, or help reflect reasoning towards the correct way.
Correct Guiding Rate (CGR)
We save reasoning processes of $FC^{1}$ samples in stage 1, and run stage 2 by including them in prompts. Similarly, we remind LLMs these reasoning may contains flaws since the samples come from the $FC^{1}$. CGR accesses whether LLMs can better reflect on their reasoning, is calculated by the changes of converting rate with and without reasoning processes.
$RPC=\frac{|TC^{1}|-|TC^{1\prime\prime}|}{|D_{VP}^{1}|},CGR=\frac{|TC^{2}|-|TC^{2\prime\prime}|}{|D_{VP}^{2}|}$
$TC^{1\prime\prime}$, $TC^{2\prime\prime}$ are obtained with previous reasoning processes included in prompts. Detailed summary of metrics is available in Table 7.
@@ 4.1 Experimental Setup
Our experiment covers both open and closed-source models: GPT-3.5 and GPT-4o (openai2023gpt4); LLaMA-3.1 series (touvron2023llama), and Qwen-2.5 series (bai2023qwen). For each model, two stimulating strategies are implemented:
@@
$\triangleright$ Concise Stimulation: Basic error notifications without correction guidance.
$\triangleright$ Detailed Stimulation: Detailed error notifications with structural and step-by-step guidance.
The experimental design considers different characteristics of two datasets forms. For Fact-based ones, which LLMs are less trained on, providing Detailed Stimulation can enhance performance; For Story-based ones, the information is embedded in long stories, which LLMs are largely trained on, providing Detailed Stimulation can induce hallucinations (Huang_2025). Experiments of different stimulation styles are in Section 5.4.
@@
To achieve a balance between randomness and dynamicity, we set all Temperatures to 0.5, except for 1.0 in Ablation Study as we want to examine the quantitative influence between prompts designing and LLMs setup.
In our results evaluation procedure, we apply a two-step output checking: the output answering is first examined for specific keywords matching. When this method fails to find one matched keyword, GPT-4o (Temperature=0.0) is deployed as a LLM judge and assign appropriate labels.
4.2 Results on the Vanilla Setting
@@
As shown in Table 3, we report detailed experimental results, including three metrics addressed before: True Converting Rate (TCR) for both stages, and Overall Converting Rate (OCR). The LLaMA and Qwen series show superior OCR on small-parameter series, whereas the GPT series shows the opposite trend. We infer that, in general, higher parameter sizes correlate with lower OCR.
4.3 Results on the RtG Setting
RtG settings contain two modes: Label and Reasoning Process (RP) Conformity: We remind LLMs that they incorrectly predicted the sample before, then add randomly assigned labels, or previous reasoning processes. A wrap-up of findings can be found in Appendix L. During our experiment, we adjust rates of assigned wrong labels at four given probabilities (100%, 66.7%, 50%, 0%). Detailed Remind-then-Guide conformity tests are available in the Table 9101112 and Appendix H.
Label Conformity
Judging from Table 3, we observe that even when LLMs are reminded of their previous incorrect answers, they still demonstrate a strong tendency towards conformity. One key influencing factor appears to be the parameter sizes: models with smaller parameter sizes, such as GPT-3.5 and LLaMA-3.1-8B, are more easily misled when reasoning solvable samples with T/F labels. In contrast, models with larger parameter sizes, such as GPT-4o and LLaMA-3.1-405B, strongly conforming to unsolvable samples with Unknown label, which is a trade-off bias regarding different sample types. Meanwhile, Qwen series show distinct yet related features, it may be attributed to smaller parameters gap between 7B and 72B compared to other model families.
Reasoning Process (RP) Conformity
We input previous reasoning processes and explicitly require LLMs to reflect on their mistakes. However, the resulting CGR values are mostly below zero, indicating that after being reminded of a past error and provided with its own flawed reasoning processes, LLMs still tend to repeat previous incorrect conclusions, leading to even larger accuracy declines.
Results on both CGR and RPC reveal that previous reasoning processes are highly unstable. Changes in accuracy can be attributed to the divergence in how problems are reasoned at different times. Notably, for unsolvable samples, reasoning processes tend to be more unreliable — reflecting on them typically leads to a higher drop in accuracy, which exposes a hidden issue: In large-scale LLM-annotated "problem-solving solutions" with own reasoning processes, unsolvable samples are particularly susceptible to unreliability. For solvable samples, the reasoning remains relatively unstable too. As a result, the precision and reliability of LLM-annotated reasoning steps for dataset samples remain to be a problem, which strongly needs manual verifications.
[…]
5 Detailed Analysis
5.1 TCR Analysis & Insights
As shown in the line chart of Figure 4, we observe clear divergences between TCR1 and TCR2 across different datasets. This indicates that many samples mis-predicted in stage 1 can be corrected in stage 2, where the model refines its initial output. In contrast, Story-based TCR1 shows a substantially lower TCR2, whereas LLaMA-3.1-8B demonstrates much higher TCR1.
Thus, we introduce model selection strategies for multi-stage reasoning workflows: (i) For GPT-4o, when facing Fact-based datasets, we can place the model at a latter order, since stage 2 offers a huge accuracy gain; However, for Story-based datasets, the improvement of stage 2 accuracy is limited, so being placed in the front should be better. (ii) For LLaMA-3.1-8B, it is always better to “do it right the first time", as being placed in the front always leads to better performance. By deploying LLMs at proper sequences, we can design better multi-stage reasoning workflows. Details are available in Appendix L.4.
5.2 Vague Perception Phenomenon Statistics
As LLMs differ in reasoning ability, we first quantify the percentage of samples that models show Vague Perception on. Bar charts in blue in Fig 4 demonstrate the percentage of samples that GPT-4o and LLaMA-3.1-8b showing this phenomenon, while complete results for all models are in Table 8.
We conclude that the Vague Perception phenomenon is prevalent across LLM Reasoning. For GPT-4o, FOLIO exhibits a relatively lower percentage. For FLD and Arts—which focus on story understanding, LLMs exhibit greater variability in understanding why samples are labeled as unknown; In sharp contrast, Science Dataset faces significant challenges even when the samples are solvable; For LLaMA-3.1-8B, there is a significant rise in the percentage of solvable samples predicted as unknown, but the ability to comprehend unknown samples is much worse.
5.3 Ablation Study
From the stage 1 output, we isolate the Unknown Converting samples ($UR^{1}$)–those LLMs have abstained on twice, and re-input them back to stage 1 stimulation. Both LLaMA-3.1-8b and GPT-4o are set with 1.0 temperature to ensure dynamicity.
As Figure 4 shows, this procedure produces virtually few converting samples: more than 95% of the samples remain unchanged. This finding demonstrates that Vague Perception is largely immune to repeated n-shot prompting; applying the same stimulation twice is therefore futile. Consequently, any effective correction must rely on a distinct stage 2 rather than reiterating stage 1, adjusting temperature alone without revising stimulation style can be limited, and all LLMs perform stably when exposed to the same prompt content multiple times.
5.4 Stimulating Styles Comparison
[…]
Furthermore, we observe a trade-off: as the reasoning ability on solvable samples declines, the model becomes better at reasoning unsolvable ones, a shift in LLM reasoning biases.
Table 4: Converting rate regarding different stimulating methods CS, DS refer to concise and detailed Stimulation. Rows shaded gray are used in our pipeline. Higher values of $TCR_{s}^{1}$ are associated with better Reliability.
Metrics(%) TCR1 $_{s}^{1} TCR$$_{u}^{1}$ Reliability
@@ 5.4 Stimulating Styles Comparison
5.5 Baseline Comparisons
In our work, we primarily analyze the “awakening ability” regarding Vague Perception samples being answered correctly based on $TC^{1}$ and $TC^{2}$ samples, and this is denoted as LLMs’ "Potential Ability". We combine this with the "Explicit Ability" Acc(Direct), as in Section 3.1, our pipeline starts with LLM reasoning on four datasets, acquiring the corresponding accuracy Acc(Direct) and extracting Vague Perception samples. By adding "Potential" and "Explicit" accuracy together, we gain the LLMs’ potential ability can be achieved without additional training or parameter revisions. Here, D represents the complete dataset, as we calculate the proportion of $TC^{1}$ and $TC^{2}$ samples over the whole dataset to acquire the accuracy. The formula is written as below:
@@ 5.5 Baseline Comparisons
$\text{Acc({WakenLLM})}=\underbrace{\text{Acc(Direct)}}_{\text{Explicit Ability}}+\underbrace{\frac{|TC^{1}\cup TC^{2}|}{|D|}}_{\text{Potential Ability}}$
As illustrated in the line chart of the left picture in Figure 5, most mainstream reasoning baselines exhibit substantial performance gaps compared to WakenLLM, indicating current reasoning frameworks are far from eliciting LLMs’ full potential, pointing to future perception-aware LLM designing. Details are in Table 6.
@@ 5.6 Phenomenon Exploration
To explore root causes of Vague Perception, we deploy GPT-4o as an LLM judge (Temperature=0.0) to identify problems based on their reasoning processes, problems mainly fall into three categories:
$\triangleright$ Fact Understanding (FU) – LLMs do not comprehend facts in the passage.
$\triangleright$ Reasoning Gap (RG) – LLMs fail to perform logical operations that reach the hypothesis.
$\triangleright$ Excessive Caution (EC) – LLMs demonstrate cautiousness over outputting their answering.
Figure 5: Reasons of samples demonstrating Vague Perception errors for GPT-4o (left). Qwen-2.5-7B and 72B False and Unexcited Converting Rate (FCR, UCR) comparison (Right). “FU”: Fact Understanding, “RG”: Reasoning Gap, and “EC”: excessive caution. Refer to caption
Causes Distribution
As shown in Figure 5, EC is a minor cause compared to FU and RG. Our experiments reveal that root causes LLMs inclined to output unknown result from both objective and subjective reasons. For the former part, FU and RG represent whether LLMs can understand given facts, or successfully inference based on comprehension, which are the main reasons; for EC, the inner biases, is a minor problem that only occurs in Story-based datasets. We attribute this to the narrative complexity of Story-based samples: long stories in Arts highlight comprehension requirements. Detailed analysis is available in Appendix J.
5.7 Degradation Analysis
In the stage 2 of WakenLLM pipeline, we don’t exclude the possibility that incorrectly predicted samples FC1 from stage 1 may be further converted to unknown in stage 2, which should be denoted as UC2. It is reasonable that LLMs initially generate incorrect answers, but because of the limited ability, they fail to output correct answers after further stimulation, and instead output unknown. We elaborate this degradation behavior as below:
As shown in Table 5, GPT-4o achieves higher TCR2 among all datasets than GPT-3.5. And problems regarding Degradation can be minor, since only GPT-4o demonstrates serious effects on Science dataset. Coupled with higher Deg, this suggests that when facing problems LLMs previously answered incorrectly—especially those beyond their ability—GPT-4o tends to output Unknown. In contrast, GPT-3.5 is much more likely to provide a definitive but incorrect answer instead of admitting uncertainty. One factor attributed to this phenomenon is LLMs’ parameter sizes.
This is further proved by the uniform distributions of UCR and FCR for Qwen-7b and 72b series. The formulas are addressed in Appendix F, calculated similarly to TCR. We compare the answering of the Qwen-2.5 series (7B and 72B model), as shown in the right radar-chart in Figure 5: (1) Models show different yet uniform output distribution among four datasets. For FLD and FOLIO, LLMs are more willing to output a false answer rather than outputting unknown; However, for Story-based datasets, like Science and Arts, LLMs behave exactly the opposite. (2) Higher the parameter sizes, the more likely models are to output unknown on Vague Perception samples.
Table 5: Degradation Analysis of GPT‑4o and GPT‑3.5. Exceptionally high Deg values of GPT-4o is in bold.
@@ 5.7 Degradation Analysis
FLD 28.19 3.08 14.43 3.09
FOLIO 44.40 5.22 42.10 6.32
Science 7.76 26.86 3.33 3.33
Arts 8.24 1.13 4.00 0.00